OpenStudy (anonymous):

Please check my answer? :) (8√6 - √2) + (5√6 - 6√2) 16√3-9 13√6 - 7√2 13√12 - 7√4 13√6 - 6√2

OpenStudy (anonymous):

so what is your answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

(my error, I thought I put it in the first half)

OpenStudy (anonymous):

I got 12 as part of the answer... 6x2

OpenStudy (anonymous):

ill do it step by step for you

OpenStudy (anonymous):

:) thanks

OpenStudy (anonymous):

\[(8\sqrt{6}-\sqrt{2})+(5\sqrt{6}-6\sqrt{2})\] \[8\sqrt{6}+5\sqrt{6}-\sqrt{2}-6\sqrt{2}\] \[(8+5)\sqrt{6}-(1+6)\sqrt{2}\] \[13\sqrt{6}-7\sqrt{2}\]

OpenStudy (anonymous):

I'm writing this down..wow thank you...!!!!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Can you see if I got this other one right please? @irkiz

OpenStudy (anonymous):

i cant get the link